OpenStudy (dragonbreath):

How many solutions does the equation have? 3(4x-2)=2(6x-3) A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

A. One solution

OpenStudy (cj49):

3(4x-2)=2(6x-3) 12x-6=12x-6 therefore infinite solutions

OpenStudy (dragonbreath):

thank you

OpenStudy (cj49):

yw mate